T) consued two watts carry ng cuuents We want hn me mos mesne a o tance from an he he SLM Do what is me of he magnehe field to E6 e) its me ma an FG compart WhAt is me magnene due ho sent nen the ware EE? what us me masnehe field due n semi-untinin were GH 2 so +OO
Answer #1

The units of r is not mentioned. I have assumed it to be cm.

Current if flowing (D to A) and (H to E)

BAB = μ0I / 2πr = (4π*10-7)(250) / (2*π*0.15) = 3.34*10-4 T (Direction: out of the plane of paper)

BCD = μ0I / 2πr = (4π*10-7)(250) / (2*π*0.15) = 3.34*10-4 T (Direction: out of the plane of paper)

Magnetic field due to circular wire at center = μ0I / 2r

Magnetic field due to quarter-circular wire at center = (1/4)(μ0I / 2r) = μ0I / 8r

BBC = μ0I / 8r = 1.05*10-6 T (Direction: out of the plane of paper)

BFG = μ0I / 8r = 1.05*10-6 T (Direction: into of the plane of paper)

BEF = BGH = 0

Net magnetic field at point P = BAB + BCD + BBC - BFG = 6.6*10-4 T (Direction: out of the plane of paper)